OpenStudy (anonymous):

One solution to the differential equation: \[y''-2\tan(x) \ y' -y=0 \ is \ y=\frac{ 1 }{ \cos x }\] Find another solution of the equation such that the two solutions form a fundamental set of solutions

OpenStudy (anonymous):

Got an A in Diff Eq in the fall, can't remember anything about it now :(

OpenStudy (anonymous):

I believe this is a problem addressing reduction of order isn't it @Wislar ? I believe you want to start with setting: \[\Large y=\frac{c(x)}{\cos x} \]

OpenStudy (anonymous):

I'm sure you could do it that way, but taking the derivatives like that would take awhile. Isn't there a way to do this with Able's formula?

OpenStudy (anonymous):

Using Abel's Formula, let \(y_2\) be the solution you're looking for. Then you set up the following equation: \[\large W(\sec x,y_2)=Ce^{-\int (-2\tan x)~dx}\] Which is the same as \[\large\begin{vmatrix}\sec x&y_2\\\sec x\tan x&y_2'\end{vmatrix}=Ce^{2\int \tan x~dx}\] Solving for \(y_2\) gives a solution in the form @Spacelimbus mentioned.

OpenStudy (anonymous):

I tried solving it using reduction of order, but I couldn't get that to work out

OpenStudy (anonymous):

\[y=\frac{ v }{ \cos x } \\ \\ y'=\frac{ v' \cos x +v \sin x }{ \cos ^{2}x } \\ y''=\frac{ v''\cos ^{3}x +vcos ^{3}x+2\cos ^{2}x \sin x v' + 2vcosx \sin ^{2}x }{ \cos ^{4}x }\]

OpenStudy (anonymous):

Have you tried the method I used? It seems to work out, and I've checked the answer with WolframAlpha

OpenStudy (anonymous):

I did try your method too and it is actually much easier than reduction of order, but I got C/cosx with yours too and that is the same thing as the current solution.

OpenStudy (anonymous):

Evaluating the Wronskian in my last equation gives you \[\large\sec x~y_2'-\sec x\tan x~y_2=Ce^{-2\ln|{\cos x}|}\\ \large\sec x~y_2'-\sec x\tan x~y_2=C\left(e^{\ln|{\cos x}|}\right)^{-2}\\ \large\sec x~y_2'-\sec x\tan x~y_2=C\left(\cos x\right)^{-2}\\ \large\sec x~y_2'-\sec x\tan x~y_2=C\sec^2x\\ \large \cos x~y_2'-\sin x~y_2=C\] Note that the LHS is the result of a differentiation of a product: \[\large \color{red}{\cos x~y_2'-\sin x~y_2}=C\\ \large \color{red}{\frac{d}{dx}\left[\cos x~y_2\right]}=C\] Integrating both sides, you get \[\large\cos x~y_2=Cx\\ \large y=\frac{Cx}{\cos x}\]

OpenStudy (anonymous):

I see what happened. I forgot to integrate the C. Thank you so much!

OpenStudy (anonymous):

You're welcome!